OpenStudy (anonymous):

y-4 = 3/4 (x-6) how do i simplify this to get a linear equation?

OpenStudy (smurfy14):

y=3x-2/4

OpenStudy (smurfy14):

put it in y=mx+b form

OpenStudy (anonymous):

what do i do with the y-4 part.?

OpenStudy (anonymous):

y-4=3/4*(x-6) add 4 to both sides y=3/4(x-6)+4 multiply right side y=(3/4)x-(3/4)*6+4 simplify y=(3/4)x-18/4+16/4 make the "4" into 16/4 (common denominator) y=(3/4)x-2/4 combine the -18/4+16/4 to get -2/4 y=(3/4)x-1/2 simplfiy [this is your answer]